\begin{thebibliography}{64}
\providecommand{\natexlab}[1]{#1}
\providecommand{\url}[1]{\texttt{#1}}
\expandafter\ifx\csname urlstyle\endcsname\relax
  \providecommand{\doi}[1]{doi: #1}\else
  \providecommand{\doi}{doi: \begingroup \urlstyle{rm}\Url}\fi

\bibitem[Acar et~al.(2021)Acar, Zhao, Matas, Mattina, Whatmough, and Saligrama]{acarfederated}
Durmus Alp~Emre Acar, Yue Zhao, Ramon Matas, Matthew Mattina, Paul Whatmough, and Venkatesh Saligrama.
\newblock {Federated Learning Based on Dynamic Regularization}.
\newblock In \emph{International Conference on Learning Representations (ICLR)}, 2021.

\bibitem[Anguita et~al.(2012)Anguita, Ghio, Oneto, Parra, and Reyes-Ortiz]{anguita2012human}
Davide Anguita, Alessandro Ghio, Luca Oneto, Xavier Parra, and Jorge~L Reyes-Ortiz.
\newblock Human activity recognition on smartphones using a multiclass hardware-friendly support vector machine.
\newblock In \emph{Ambient Assisted Living and Home Care: 4th International Workshop, IWAAL 2012, Vitoria-Gasteiz, Spain, December 3-5, 2012. Proceedings 4}, pages 216--223. Springer, 2012.

\bibitem[Arivazhagan et~al.(2019)Arivazhagan, Aggarwal, Singh, and Choudhary]{arivazhagan2019federated}
Manoj~Ghuhan Arivazhagan, Vinay Aggarwal, Aaditya~Kumar Singh, and Sunav Choudhary.
\newblock {Federated Learning with Personalization Layers}.
\newblock \emph{arXiv preprint arXiv:1912.00818}, 2019.

\bibitem[Asad et~al.(2020)Asad, Moustafa, and Ito]{asad2020fedopt}
Muhammad Asad, Ahmed Moustafa, and Takayuki Ito.
\newblock Fedopt: Towards communication efficiency and privacy preservation in federated learning.
\newblock \emph{Applied Sciences}, 10\penalty0 (8):\penalty0 2864, 2020.

\bibitem[Beutel et~al.(2020)Beutel, Topal, Mathur, Qiu, Fernandez-Marques, Gao, Sani, Li, Parcollet, de~Gusm{\~a}o, et~al.]{beutel2020flower}
Daniel~J Beutel, Taner Topal, Akhil Mathur, Xinchi Qiu, Javier Fernandez-Marques, Yan Gao, Lorenzo Sani, Kwing~Hei Li, Titouan Parcollet, Pedro Porto~Buarque de~Gusm{\~a}o, et~al.
\newblock Flower: A friendly federated learning research framework.
\newblock \emph{arXiv preprint arXiv:2007.14390}, 2020.

\bibitem[Caldas et~al.(2018)Caldas, Duddu, Wu, Li, Kone{\v{c}}n{\`y}, McMahan, Smith, and Talwalkar]{caldas2018leaf}
Sebastian Caldas, Sai Meher~Karthik Duddu, Peter Wu, Tian Li, Jakub Kone{\v{c}}n{\`y}, H~Brendan McMahan, Virginia Smith, and Ameet Talwalkar.
\newblock Leaf: A benchmark for federated settings.
\newblock \emph{arXiv preprint arXiv:1812.01097}, 2018.

\bibitem[Chai et~al.(2020)Chai, Wang, Yang, Zhang, Chen, and Yang]{chai2020fedeval}
Di~Chai, Leye Wang, Liu Yang, Junxue Zhang, Kai Chen, and Qiang Yang.
\newblock Fedeval: A holistic evaluation framework for federated learning.
\newblock \emph{arXiv preprint arXiv:2011.09655}, 2020.

\bibitem[Chen et~al.(2022)Chen, Gao, Kuang, Li, and Ding]{chen2022pfl}
Daoyuan Chen, Dawei Gao, Weirui Kuang, Yaliang Li, and Bolin Ding.
\newblock pfl-bench: A comprehensive benchmark for personalized federated learning.
\newblock \emph{Advances in Neural Information Processing Systems}, 35:\penalty0 9344--9360, 2022.

\bibitem[Chen and Chao(2021)]{chen2021bridging}
Hong-You Chen and Wei-Lun Chao.
\newblock {On Bridging Generic and Personalized Federated Learning for Image Classification}.
\newblock In \emph{International Conference on Learning Representations (ICLR)}, 2021.

\bibitem[Chrabaszcz et~al.(2017)Chrabaszcz, Loshchilov, and Hutter]{chrabaszcz2017downsampled}
Patryk Chrabaszcz, Ilya Loshchilov, and Frank Hutter.
\newblock {A Downsampled Variant of Imagenet as an Alternative to the Cifar Datasets}.
\newblock \emph{arXiv preprint arXiv:1707.08819}, 2017.

\bibitem[Collins et~al.(2021)Collins, Hassani, Mokhtari, and Shakkottai]{collins2021exploiting}
Liam Collins, Hamed Hassani, Aryan Mokhtari, and Sanjay Shakkottai.
\newblock {Exploiting Shared Representations for Personalized Federated Learning}.
\newblock In \emph{International Conference on Machine Learning (ICML)}, 2021.

\bibitem[Deng et~al.(2020)Deng, Kamani, and Mahdavi]{deng2020adaptive}
Yuyang Deng, Mohammad~Mahdi Kamani, and Mehrdad Mahdavi.
\newblock {Adaptive Personalized Federated Learning}.
\newblock \emph{arXiv preprint arXiv:2003.13461}, 2020.

\bibitem[Fallah et~al.(2020)Fallah, Mokhtari, and Ozdaglar]{NEURIPS2020_24389bfe}
Alireza Fallah, Aryan Mokhtari, and Asuman Ozdaglar.
\newblock {Personalized Federated Learning with Theoretical Guarantees: A Model-Agnostic Meta-Learning Approach}.
\newblock In \emph{International Conference on Advances in Neural Information Processing Systems (NeurIPS)}, 2020.

\bibitem[He et~al.(2020)He, Li, So, Zeng, Zhang, Wang, Wang, Vepakomma, Singh, Qiu, et~al.]{he2020fedml}
Chaoyang He, Songze Li, Jinhyun So, Xiao Zeng, Mi~Zhang, Hongyi Wang, Xiaoyang Wang, Praneeth Vepakomma, Abhishek Singh, Hang Qiu, et~al.
\newblock Fedml: A research library and benchmark for federated machine learning.
\newblock \emph{arXiv preprint arXiv:2007.13518}, 2020.

\bibitem[He et~al.(2016)He, Zhang, Ren, and Sun]{he2016deep}
Kaiming He, Xiangyu Zhang, Shaoqing Ren, and Jian Sun.
\newblock {Deep Residual Learning for Image Recognition}.
\newblock In \emph{IEEE Conference on Computer Vision and Pattern Recognition (CVPR)}, 2016.

\bibitem[Hu et~al.(2022)Hu, Li, Liu, Li, Wu, and He]{hu2022oarf}
Sixu Hu, Yuan Li, Xu~Liu, Qinbin Li, Zhaomin Wu, and Bingsheng He.
\newblock The oarf benchmark suite: Characterization and implications for federated learning systems.
\newblock \emph{ACM Transactions on Intelligent Systems and Technology (TIST)}, 13\penalty0 (4):\penalty0 1--32, 2022.

\bibitem[Huang et~al.(2021)Huang, Chu, Zhou, Wang, Liu, Pei, and Zhang]{huang2021personalized}
Yutao Huang, Lingyang Chu, Zirui Zhou, Lanjun Wang, Jiangchuan Liu, Jian Pei, and Yong Zhang.
\newblock {Personalized Cross-Silo Federated Learning on Non-IID Data}.
\newblock In \emph{AAAI Conference on Artificial Intelligence (AAAI)}, 2021.

\bibitem[Joulin et~al.(2017)Joulin, Grave, Bojanowski, and Mikolov]{joulinetal2017bag}
Armand Joulin, Edouard Grave, Piotr Bojanowski, and Tomas Mikolov.
\newblock {Bag of Tricks for Efficient Text Classification}.
\newblock In \emph{Conference of the European Chapter of the Association for Computational Linguistics (EACL)}, 2017.

\bibitem[Karimireddy et~al.(2020)Karimireddy, Kale, Mohri, Reddi, Stich, and Suresh]{karimireddy2020scaffold}
Sai~Praneeth Karimireddy, Satyen Kale, Mehryar Mohri, Sashank Reddi, Sebastian Stich, and Ananda~Theertha Suresh.
\newblock {Scaffold: Stochastic Controlled Averaging for Federated Learning}.
\newblock In \emph{International Conference on Machine Learning (ICML)}, 2020.

\bibitem[Koh et~al.(2021)Koh, Sagawa, Marklund, Xie, Zhang, Balsubramani, Hu, Yasunaga, Phillips, Gao, et~al.]{koh2021wilds}
Pang~Wei Koh, Shiori Sagawa, Henrik Marklund, Sang~Michael Xie, Marvin Zhang, Akshay Balsubramani, Weihua Hu, Michihiro Yasunaga, Richard~Lanas Phillips, Irena Gao, et~al.
\newblock Wilds: A benchmark of in-the-wild distribution shifts.
\newblock In \emph{International conference on machine learning}, 2021.

\bibitem[LeCun et~al.(1998)LeCun, Bottou, Bengio, and Haffner]{lecun1998gradient}
Yann LeCun, L{\'e}on Bottou, Yoshua Bengio, and Patrick Haffner.
\newblock {Gradient-Based Learning Applied to Document Recognition}.
\newblock \emph{Proceedings of the IEEE}, 86\penalty0 (11):\penalty0 2278--2324, 1998.

\bibitem[Lee et~al.(2022)Lee, Jeong, Shin, Bae, and Yun]{lee2022preservation}
Gihun Lee, Minchan Jeong, Yongjin Shin, Sangmin Bae, and Se-Young Yun.
\newblock Preservation of the global knowledge by not-true distillation in federated learning.
\newblock \emph{Advances in Neural Information Processing Systems}, 35:\penalty0 38461--38474, 2022.

\bibitem[Li et~al.(2021{\natexlab{a}})Li, He, and Song]{li2021model}
Qinbin Li, Bingsheng He, and Dawn Song.
\newblock {Model-Contrastive Federated Learning}.
\newblock In \emph{IEEE Conference on Computer Vision and Pattern Recognition (CVPR)}, 2021{\natexlab{a}}.

\bibitem[Li et~al.(2022)Li, Diao, Chen, and He]{li2022federated}
Qinbin Li, Yiqun Diao, Quan Chen, and Bingsheng He.
\newblock Federated learning on non-iid data silos: An experimental study.
\newblock In \emph{2022 IEEE 38th international conference on data engineering (ICDE)}, pages 965--978. IEEE, 2022.

\bibitem[Li et~al.(2020)Li, Sahu, Zaheer, Sanjabi, Talwalkar, and Smith]{MLSYS2020_38af8613}
Tian Li, Anit~Kumar Sahu, Manzil Zaheer, Maziar Sanjabi, Ameet Talwalkar, and Virginia Smith.
\newblock {Federated Optimization in Heterogeneous Networks}.
\newblock In \emph{Conference on Machine Learning and Systems (MLSys)}, 2020.

\bibitem[Li et~al.(2021{\natexlab{b}})Li, Hu, Beirami, and Smith]{li2021ditto}
Tian Li, Shengyuan Hu, Ahmad Beirami, and Virginia Smith.
\newblock {Ditto: Fair and Robust Federated Learning Through Personalization}.
\newblock In \emph{International Conference on Machine Learning (ICML)}, 2021{\natexlab{b}}.

\bibitem[Li et~al.(2021{\natexlab{c}})Li, JIANG, Zhang, Kamp, and Dou]{li2020fedbn}
Xiaoxiao Li, Meirui JIANG, Xiaofei Zhang, Michael Kamp, and Qi~Dou.
\newblock {FedBN: Federated Learning on Non-IID Features via Local Batch Normalization}.
\newblock In \emph{International Conference on Learning Representations (ICLR)}, 2021{\natexlab{c}}.

\bibitem[Li et~al.(2021{\natexlab{d}})Li, Zhan, Shao, Li, and Song]{li2021fedphp}
Xin-Chun Li, De-Chuan Zhan, Yunfeng Shao, Bingshuai Li, and Shaoming Song.
\newblock {FedPHP: Federated Personalization with Inherited Private Models}.
\newblock In \emph{European Conference on Machine Learning and Principles and Practice of Knowledge Discovery in Databases (ECML)}, 2021{\natexlab{d}}.

\bibitem[Liang et~al.(2020)Liang, Liu, Ziyin, Allen, Auerbach, Brent, Salakhutdinov, and Morency]{liang2020think}
Paul~Pu Liang, Terrance Liu, Liu Ziyin, Nicholas~B Allen, Randy~P Auerbach, David Brent, Ruslan Salakhutdinov, and Louis-Philippe Morency.
\newblock Think locally, act globally: Federated learning with local and global representations.
\newblock \emph{arXiv preprint arXiv:2001.01523}, 2020.

\bibitem[Liu et~al.(2021)Liu, Fan, Chen, Xu, and Yang]{liu2021fate}
Yang Liu, Tao Fan, Tianjian Chen, Qian Xu, and Qiang Yang.
\newblock Fate: An industrial grade platform for collaborative learning with data protection.
\newblock \emph{The Journal of Machine Learning Research}, 22\penalty0 (1):\penalty0 10320--10325, 2021.

\bibitem[Luo and Wu(2022)]{ijcai2022p301}
Jun Luo and Shandong Wu.
\newblock {Adapt to Adaptation: Learning Personalization for Cross-Silo Federated Learning}.
\newblock In \emph{International Joint Conference on Artificial Intelligence (IJCAI)}, 2022.

\bibitem[McMahan et~al.(2017)McMahan, Moore, Ramage, Hampson, and y~Arcas]{mcmahan2017communication}
Brendan McMahan, Eider Moore, Daniel Ramage, Seth Hampson, and Blaise~Aguera y~Arcas.
\newblock {Communication-Efficient Learning of Deep Networks from Decentralized Data}.
\newblock In \emph{International Conference on Artificial Intelligence and Statistics (AISTATS)}, 2017.

\bibitem[Niu and Deng(2022)]{niu2022federated}
Yifan Niu and Weihong Deng.
\newblock {Federated Learning for Face Recognition With Gradient Correction}.
\newblock In \emph{AAAI Conference on Artificial Intelligence (AAAI)}, 2022.

\bibitem[Oh et~al.(2022)Oh, Kim, and Yun]{ohfedbabu}
Jaehoon Oh, SangMook Kim, and Se-Young Yun.
\newblock {FedBABU: Toward Enhanced Representation for Federated Image Classification}.
\newblock In \emph{International Conference on Learning Representations (ICLR)}, 2022.

\bibitem[Reiss and Stricker(2012)]{reiss2012introducing}
Attila Reiss and Didier Stricker.
\newblock Introducing a new benchmarked dataset for activity monitoring.
\newblock In \emph{2012 16th international symposium on wearable computers}, pages 108--109. IEEE, 2012.

\bibitem[Seo et~al.(2022)Seo, Park, Oh, Bennis, and Kim]{seo202216}
Hyowoon Seo, Jihong Park, Seungeun Oh, Mehdi Bennis, and Seong-Lyun Kim.
\newblock 16 federated knowledge distillation.
\newblock \emph{Machine Learning and Wireless Communications}, page 457, 2022.

\bibitem[Shen et~al.(2020)Shen, Zhang, Jia, Zhang, Huang, Zhou, Kuang, Wu, and Wu]{shen2020federated}
Tao Shen, Jie Zhang, Xinkang Jia, Fengda Zhang, Gang Huang, Pan Zhou, Kun Kuang, Fei Wu, and Chao Wu.
\newblock Federated mutual learning.
\newblock \emph{arXiv preprint arXiv:2006.16765}, 2020.

\bibitem[Shi and Kontar(2023)]{shi2023personalized}
Naichen Shi and Raed~Al Kontar.
\newblock Personalized federated learning via domain adaptation with an application to distributed 3d printing.
\newblock \emph{Technometrics}, 65\penalty0 (3):\penalty0 328--339, 2023.

\bibitem[T~Dinh et~al.(2020)T~Dinh, Tran, and Nguyen]{t2020personalized}
Canh T~Dinh, Nguyen Tran, and Tuan~Dung Nguyen.
\newblock {Personalized Federated Learning with Moreau Envelopes}.
\newblock In \emph{International Conference on Advances in Neural Information Processing Systems (NeurIPS)}, 2020.

\bibitem[Tan et~al.(2022{\natexlab{a}})Tan, Yu, Cui, and Yang]{tan2022towards}
Alysa~Ziying Tan, Han Yu, Lizhen Cui, and Qiang Yang.
\newblock {Towards Personalized Federated Learning}.
\newblock \emph{IEEE Transactions on Neural Networks and Learning Systems}, 2022{\natexlab{a}}.
\newblock Early Access.

\bibitem[Tan et~al.(2022{\natexlab{b}})Tan, Long, Liu, Zhou, Lu, Jiang, and Zhang]{tan2022fedproto}
Yue Tan, Guodong Long, Lu~Liu, Tianyi Zhou, Qinghua Lu, Jing Jiang, and Chengqi Zhang.
\newblock {Fedproto: Federated Prototype Learning across Heterogeneous Clients}.
\newblock In \emph{AAAI Conference on Artificial Intelligence (AAAI)}, 2022{\natexlab{b}}.

\bibitem[Tan et~al.(2022{\natexlab{c}})Tan, Long, Ma, Liu, Zhou, and Jiang]{tan2022federated}
Yue Tan, Guodong Long, Jie Ma, Lu~Liu, Tianyi Zhou, and Jing Jiang.
\newblock Federated learning from pre-trained models: A contrastive learning approach.
\newblock \emph{arXiv preprint arXiv:2209.10083}, 2022{\natexlab{c}}.

\bibitem[Wu et~al.(2022{\natexlab{a}})Wu, Wu, Lyu, Huang, and Xie]{wu2022communication}
Chuhan Wu, Fangzhao Wu, Lingjuan Lyu, Yongfeng Huang, and Xing Xie.
\newblock Communication-efficient federated learning via knowledge distillation.
\newblock \emph{Nature communications}, 13\penalty0 (1):\penalty0 2032, 2022{\natexlab{a}}.

\bibitem[Wu et~al.(2022{\natexlab{b}})Wu, Li, Charles, Xiao, Liu, Xu, and Smith]{wu2022motley}
Shanshan Wu, Tian Li, Zachary Charles, Yu~Xiao, Ken Liu, Zheng Xu, and Virginia Smith.
\newblock Motley: Benchmarking heterogeneity and personalization in federated learning.
\newblock In \emph{Workshop on Federated Learning: Recent Advances and New Challenges (in Conjunction with NeurIPS 2022)}, 2022{\natexlab{b}}.

\bibitem[Wu et~al.(2023)Wu, Liu, Niu, Zhu, and Tang]{wu2023bold}
Xinghao Wu, Xuefeng Liu, Jianwei Niu, Guogang Zhu, and Shaojie Tang.
\newblock Bold but cautious: Unlocking the potential of personalized federated learning through cautiously aggressive collaboration.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on Computer Vision}, 2023.

\bibitem[Wu et~al.(2022{\natexlab{c}})Wu, Kang, Luo, He, Fan, Pan, and Yang]{wufedcg}
Yuezhou Wu, Yan Kang, Jiahuan Luo, Yuanqin He, Lixin Fan, Rong Pan, and Qiang Yang.
\newblock Fedcg: Leverage conditional gan for protecting privacy and maintaining competitive performance in federated learning.
\newblock In \emph{International Conference on Advances in Neural Information Processing Systems (NeurIPS)}, 2022{\natexlab{c}}.

\bibitem[Xiao et~al.(2017)Xiao, Rasul, and Vollgraf]{xiao2017fashion}
Han Xiao, Kashif Rasul, and Roland Vollgraf.
\newblock {Fashion-MNIST: A Novel Image Dataset for Benchmarking Machine Learning Algorithms}.
\newblock \emph{arXiv preprint arXiv:1708.07747}, 2017.

\bibitem[Xie et~al.(2023)Xie, Wang, Gao, Chen, Yao, Kuang, Li, Ding, and Zhou]{federatedscope}
Yuexiang Xie, Zhen Wang, Dawei Gao, Daoyuan Chen, Liuyi Yao, Weirui Kuang, Yaliang Li, Bolin Ding, and Jingren Zhou.
\newblock Federatedscope: A flexible federated learning platform for heterogeneity.
\newblock \emph{Proceedings of the VLDB Endowment}, 16\penalty0 (5):\penalty0 1059--1072, 2023.

\bibitem[Xu et~al.(2022)Xu, Tong, and Huang]{xu2022personalized}
Jian Xu, Xinyi Tong, and Shao-Lun Huang.
\newblock Personalized federated learning with feature alignment and classifier collaboration.
\newblock In \emph{The Eleventh International Conference on Learning Representations}, 2022.

\bibitem[Yang et~al.(2019)Yang, Liu, Chen, and Tong]{yang2019federated}
Qiang Yang, Yang Liu, Tianjian Chen, and Yongxin Tong.
\newblock {Federated Machine Learning: Concept and Applications}.
\newblock \emph{ACM Transactions on Intelligent Systems and Technology}, 10\penalty0 (2):\penalty0 1--19, 2019.

\bibitem[Yi et~al.(2023)Yi, Wang, Liu, Shi, and Yu]{yi2023fedgh}
Liping Yi, Gang Wang, Xiaoguang Liu, Zhuan Shi, and Han Yu.
\newblock Fedgh: Heterogeneous federated learning with generalized global header.
\newblock In \emph{Proceedings of the 31st ACM International Conference on Multimedia}, 2023.

\bibitem[Zeng et~al.(2023)Zeng, Liang, Hu, Wang, and Xu]{zeng2023fedlab}
Dun Zeng, Siqi Liang, Xiangjing Hu, Hui Wang, and Zenglin Xu.
\newblock Fedlab: A flexible federated learning framework.
\newblock \emph{J. Mach. Learn. Res.}, 24:\penalty0 100--1, 2023.

\bibitem[Zhang et~al.(2023{\natexlab{a}})Zhang, Hua, Cao, Wang, Song, Xue, Ma, and Guan]{zhang2023eliminating}
Jianqing Zhang, Yang Hua, Jian Cao, Hao Wang, Tao Song, Zhengui Xue, Ruhui Ma, and Haibing Guan.
\newblock Eliminating domain bias for federated learning in representation space.
\newblock \emph{Advances in Neural Information Processing Systems}, 36, 2023{\natexlab{a}}.

\bibitem[Zhang et~al.(2023{\natexlab{b}})Zhang, Hua, Wang, Song, Xue, Ma, Cao, and Guan]{zhang2023gpfl}
Jianqing Zhang, Yang Hua, Hao Wang, Tao Song, Zhengui Xue, Ruhui Ma, Jian Cao, and Haibing Guan.
\newblock Gpfl: Simultaneously learning global and personalized feature information for personalized federated learning.
\newblock In \emph{Proceedings of the IEEE/CVF International Conference on Computer Vision}, pages 5041--5051, 2023{\natexlab{b}}.

\bibitem[Zhang et~al.(2023{\natexlab{c}})Zhang, Hua, Wang, Song, Xue, Ma, and Guan]{Zhang2023fedcp}
Jianqing Zhang, Yang Hua, Hao Wang, Tao Song, Zhengui Xue, Ruhui Ma, and Haibing Guan.
\newblock Fedcp: Separating feature information for personalized federated learning via conditional policy.
\newblock In \emph{Proceedings of the 29th ACM SIGKDD Conference on Knowledge Discovery and Data Mining}, 2023{\natexlab{c}}.

\bibitem[Zhang et~al.(2023{\natexlab{d}})Zhang, Hua, Wang, Song, Xue, Ma, and Guan]{zhang2023fedala}
Jianqing Zhang, Yang Hua, Hao Wang, Tao Song, Zhengui Xue, Ruhui Ma, and Haibing Guan.
\newblock {FedALA: Adaptive Local Aggregation for Personalized Federated Learning}.
\newblock In \emph{AAAI Conference on Artificial Intelligence (AAAI)}, 2023{\natexlab{d}}.

\bibitem[Zhang et~al.(2024{\natexlab{a}})Zhang, Liu, Hua, and Cao]{zhang2024fedtgp}
Jianqing Zhang, Yang Liu, Yang Hua, and Jian Cao.
\newblock Fedtgp: Trainable global prototypes with adaptive-margin-enhanced contrastive learning for data and model heterogeneity in federated learning.
\newblock In \emph{Proceedings of the AAAI Conference on Artificial Intelligence}, 2024{\natexlab{a}}.

\bibitem[Zhang et~al.(2024{\natexlab{b}})Zhang, Liu, Hua, and Cao]{zhang2024upload}
Jianqing Zhang, Yang Liu, Yang Hua, and Jian Cao.
\newblock An upload-efficient scheme for transferring knowledge from a server-side pre-trained generator to clients in heterogeneous federated learning.
\newblock In \emph{Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition}, 2024{\natexlab{b}}.

\bibitem[Zhang et~al.(2024{\natexlab{c}})Zhang, Liu, Hua, Cao, and Yang]{zhang2024fedl2g}
Jianqing Zhang, Yang Liu, Yang Hua, Jian Cao, and Qiang Yang.
\newblock Fedl2g: Learning to guide local training in heterogeneous federated learning.
\newblock \emph{arXiv preprint arXiv:2410.06490}, 2024{\natexlab{c}}.

\bibitem[Zhang et~al.(2022)Zhang, Li, Li, Xu, Wu, Ding, and Wu]{zhang2022federated1}
Jie Zhang, Zhiqi Li, Bo~Li, Jianghe Xu, Shuang Wu, Shouhong Ding, and Chao Wu.
\newblock Federated learning with label distribution skew via logits calibration.
\newblock In \emph{International Conference on Machine Learning}, 2022.

\bibitem[Zhang et~al.(2020)Zhang, Sapra, Fidler, Yeung, and Alvarez]{zhang2020personalized}
Michael Zhang, Karan Sapra, Sanja Fidler, Serena Yeung, and Jose~M Alvarez.
\newblock {Personalized Federated Learning with First Order Model Optimization}.
\newblock In \emph{International Conference on Learning Representations (ICLR)}, 2020.

\bibitem[Zhang et~al.(2015)Zhang, Zhao, and LeCun]{zhang2015character}
Xiang Zhang, Junbo Zhao, and Yann LeCun.
\newblock {Character-Level Convolutional Networks for Text Classification}.
\newblock In \emph{International Conference on Advances in Neural Information Processing Systems (NeurIPS)}, 2015.

\bibitem[Zhu et~al.(2019)Zhu, Liu, and Han]{zhu2019deep}
Ligeng Zhu, Zhijian Liu, and Song Han.
\newblock {Deep Leakage from Gradients}.
\newblock In \emph{International Conference on Advances in Neural Information Processing Systems (NeurIPS)}, 2019.

\bibitem[Zhu et~al.(2021)Zhu, Hong, and Zhou]{zhu2021data}
Zhuangdi Zhu, Junyuan Hong, and Jiayu Zhou.
\newblock {Data-Free Knowledge Distillation for Heterogeneous Federated Learning}.
\newblock In \emph{International Conference on Machine Learning (ICML)}, 2021.

\end{thebibliography}

\documentclass[twoside,11pt]{article}

\usepackage{blindtext}

\usepackage{jmlr2e}
\usepackage{xcolor}         % colors

\usepackage{lastpage}
\jmlrheading{26}{2025}{1-\pageref{LastPage}}{12/23}{2/25}{23-1634}{Jianqing Zhang, Yang Liu, Yang Hua, Hao Wang, Tao Song, Zhengui Xue, Ruhui Ma, and Jian Cao}

\ShortHeadings{\pfllib: A Personalized Federated Learning Library and Benchmark Platform}{Zhang, Liu, Hua, Wang, Song, Xue, Ma, and Cao}
\firstpageno{1}

\makeatletter
\DeclareRobustCommand\onedot{\futurelet\@let@token\@onedot}
\def\@onedot{\ifx\@let@token.\else.\null\fi\xspace}

\makeatother

\usepackage[capitalize]{cleveref}
\crefname{section}{Sec.}{Secs.}
\Crefname{section}{Section}{Sections}
\Crefname{table}{Table}{Tables}
\crefname{table}{Tab.}{Tabs.}

\usepackage{subcaption}
\usepackage{booktabs}

\usepackage{xspace}
\def\pfllib{\texttt{PFLlib}\xspace}

\definecolor{SJTUred}{RGB}{202, 21, 30}
\newcommand{\linecode}[1]{\colorbox[rgb]{1,1,1}{\color{SJTUred} \texttt{#1}}}

\usepackage{listings}
\usepackage{xcolor}
\definecolor{codegreen}{rgb}{0,0.6,0}
\definecolor{codegray}{rgb}{0.5,0.5,0.5}
\definecolor{codeblue}{RGB}{88,160,211}
\definecolor{backcolour}{RGB}{238,238,238}
\lstdefinestyle{mystyle}{
    backgroundcolor=\color{backcolour},   
    commentstyle=\color{codegreen},
    keywordstyle=\color{codeblue},
    numberstyle=\tiny\color{codegray},
    stringstyle=\color{codegreen},
    basicstyle=\ttfamily\footnotesize,
    breakatwhitespace=false,         
    breaklines=true,                 
    captionpos=b,                    
    keepspaces=true,                 
    numbers=left,                    
    numbersep=5pt,                  
    showspaces=false,                
    showstringspaces=false,
    showtabs=false,                  
    tabsize=2
}
\usepackage{multirow}
\usepackage{array}

\hypersetup{hidelinks}

\begin{document}

\title{\pfllib: A Beginner-Friendly and Comprehensive Personalized Federated Learning Library and Benchmark}

\author{\name Jianqing Zhang$^1$ \email tsingz@sjtu.edu.cn \\
\name Yang Liu$^{2,*}$ \email liuy03@air.tsinghua.edu.cn \\
\name Yang Hua$^3$ \email y.hua@qub.ac.uk \\
\name Hao Wang$^4$ \email hwang9@stevens.edu \\
\name Tao Song$^1$ \email songt333@sjtu.edu.cn \\
\name Zhengui Xue$^1$ \email zhenguixue@sjtu.edu.cn \\
\name Ruhui Ma$^1$ \email ruhuima@sjtu.edu.cn \\
\name Jian Cao$^{1,*}$ \email cao-jian@sjtu.edu.cn \\
\addr $^1$ Shanghai Jiao Tong University, Shanghai, China\\
\addr $^2$ Institute for AI Industry Research, Tsinghua University, Beijing, China\\
\addr $^3$ Queen's University Belfast, Belfast, UK\\
\addr $^4$ Stevens Institute of Technology, New Jersey, USA\\
\addr $^*$ Corresponding Authors
}

\editor{Albert Bifet}

\maketitle

\begin{abstract}%   <- trailing '%' for backward compatibility of .sty file
Amid the ongoing advancements in Federated Learning (FL), a machine learning paradigm that allows collaborative learning with data privacy protection, personalized FL (pFL) has gained significant prominence as a research direction within the FL domain. Whereas traditional FL (tFL) focuses on jointly learning a global model, pFL aims to balance each client's global and personalized goals in FL settings. To foster the pFL research community, we started and built \pfllib, a comprehensive pFL library with an integrated benchmark platform. In \pfllib, we implemented 37 state-of-the-art FL algorithms (8 tFL algorithms and 29 pFL algorithms) and provided various evaluation environments with three statistically heterogeneous scenarios and 24 datasets. 
At present, \pfllib\footnote{\url{https://www.pfllib.com/} and \url{https://github.com/TsingZ0/PFLlib}} has gained more than 1600 stars and 300 forks on GitHub. 
\end{abstract}

\begin{keywords}
  federated learning, personalization, privacy, benchmark, heterogeneity
\end{keywords}

\section{Introduction}

Federated Learning (FL) has gained significant attention due to its ability to perform distributed machine learning while ensuring privacy preservation~\citep{yang2019federated}. In traditional FL (tFL) algorithms, such as FedAvg~\citep{mcmahan2017communication}, participating clients train local models using local data and send only local model updates to a global server, which then aggregates these updates to obtain a global model. These approaches do not consider the customization needs of each local client. %However, in practice, the Artificial Intelligence (AI) model trained collaboratively using traditional FL (tFL) algorithms, such as FedAvg~\citep{mcmahan2017communication}, faces the challenge of statistical heterogeneity. This includes issues such as not non-independent and identically distributed (Non-IID) data among the participating clients. Next, 
Personalized FL (pFL) is introduced to train customized client models to improve their performance on individualized tasks. %unlike tFL approaches that only train a single global model. 
In tandem with the burgeoning prominence of pFL, there has been a surge in the development of various pFL algorithms and associated techniques~\citep{tan2022towards, zhang2023gpfl}. However, due to their rapid progress and diverse settings, the difficulties of tracking, implementing, and benchmarking these methods also grow tremendously. 

To alleviate these challenges, we have developed \pfllib, a comprehensive pFL library with an integrated benchmark platform. \pfllib includes implementations of \textbf{37} state-of-the-art (SOTA) tFL and pFL algorithms, encompassing 8 tFL algorithms and \textbf{29} pFL algorithms. Our library is beginner-friendly and easily extensible, allowing contributors to seamlessly add new algorithms, scenarios, and datasets, thus ensuring that \pfllib remains up-to-date and popular. In addition, we have implemented three types of data heterogeneity scenarios and incorporated 24 datasets, covering Computer Vision (CV), Natural Language Processing (NLP), and Sensor Signal Processing (SSP) tasks. We can evaluate FL algorithms in \pfllib and assess their adaptability to various scenarios, providing valuable information for algorithm selection and evaluation in practical applications. 

\section{Related Work}

With the rapid development of the FL field, numerous benchmarks and platforms have emerged in recent years. Most of their latest versions are for practical deployments, such as FATE~\citep{liu2021fate}, FedML~\citep{he2020fedml}, 
FederatedScope~\citep{federatedscope}, Flower~\citep{beutel2020flower}, TensorFlow Federated\footnote{\url{https://www.tensorflow.org/federated}}, NVIDIA Clara\footnote{\url{https://developer.nvidia.com/industries/healthcare}}, SecretFlow\footnote{\url{https://github.com/secretflow/secretflow}}, Fedlearner\footnote{\url{https://github.com/bytedance/fedlearner}}, and
PySyft\footnote{\url{https://github.com/OpenMined/PySyft}}. 
% \yang{many of these projects are not "built for industrial applications" only, I would argue, such as FedML, PySyft, Tensorflow TF. I think it is better to stress that existing projects lack coverage on pFL algorithms, followed by a detailed discussion on pFL coverage in existing FL projects. I.e., are there pFL projects? How to compare ours against them in terms of coverage, implementation etc.}
Despite the efficient resource management and extensive functionality offered by these platforms, they can present a challenge for beginners who seek to comprehend the fundamental mechanisms of FL and delve into the philosophical aspects of existing FL algorithms. There are also some beginner-friendly platforms, such as 
LEAF~\citep{caldas2018leaf}, NIID-Bench~\citep{li2022federated}, Motley~\citep{wu2022motley}, OARF~\citep{hu2022oarf}, FedEval~\citep{chai2020fedeval}, and FedLab~\citep{zeng2023fedlab}.
However, these benchmarks and platforms still lack sufficient and up-to-date built-in SOTA pFL algorithms for researchers to learn, compare, and analyze. 
% \yang{need to be more detailed in comparison. I remember we discussed some recent PFL projects which are accepted by major conferences, are they all included?}

pFL-Bench~\citep{chen2022pfl} is one of the latest PFL projects, %which includes 20 pFL methods. However, there are 
but it supports only 5 SOTA pFL methods while the remaining pFL methods are variants created by combining them with existing approaches including FedBN~\citep{li2020fedbn}, FedOpt~\citep{asad2020fedopt}, and Fine-tuning (FT). Besides, all the pFL algorithms in pFL-Bench are outdated (before 2022). 
% Besides, it is also difficult for pFL-Bench to classify pFL methods with limited SOTA pFL methods. 
In contrast, our \pfllib consists of 29 SOTA pFL algorithms. Moreover, due to our straightforward file structure, \pfllib is more accessible for beginners to learn and utilize pFL algorithms compared to the complex pFL-Bench. 
% \yang{any other differences/advantages of ours compared to pFL-Bench worth mentioning? In terms of how to use, extend, and evaluate maybe? e.g., metrics used to evaluate, efficiency, privacy etc. }
% We also classify these 29 pFL algorithms into 6 categories. 

\section{\pfllib: A Beginner-Friendly and Comprehensive Library}

\noindent\textbf{Algorithms.} 
In our \pfllib, the primary focus is on pFL algorithms. In addition, we have also included a selection of tFL algorithms to facilitate the evaluation of pFL algorithms, following previous pFL research ~\citep{li2021ditto, t2020personalized, zhang2023fedala}. Based on their foundational techniques, we have categorized 8 tFL algorithms and 29 pFL algorithms. The detailed classification is presented in \Cref{tab:algo}. 

\begin{table}[h]
  \centering
  \caption{FL Algorithm Taxonomy in our \pfllib.}
  \resizebox{\linewidth}{!}{
    \begin{tabular}{l|l|p{105mm}}
    \toprule
     & Category & Algorithms \\
    \midrule
    \parbox[t]{2mm}{\multirow{6}{*}{\rotatebox[origin=c]{90}{8 tFL Algorithms}}} & Basic tFL & FedAvg~\citep{mcmahan2017communication} \\
    \cmidrule{2-3}
    & Update-correction-based tFL & SCAFFOLD~\citep{karimireddy2020scaffold} \\
    \cmidrule{2-3}
    & Regularization-based tFL & FedProx~\citep{MLSYS2020_38af8613} and FedDyn~\citep{acarfederated} \\
    \cmidrule{2-3}
    & Model-splitting-based tFL & MOON~\citep{li2021model} and FedLC~\citep{zhang2022federated1} \\
    \cmidrule{2-3}
    & Knowledge-distillation-based tFL & FedGen~\citep{zhu2021data} and FedNTD~\citep{lee2022preservation} \\
    \midrule
    \parbox[t]{2mm}{\multirow{15}{*}{\rotatebox[origin=c]{90}{29 pFL Algorithms}}} & Meta-learning-based pFL & Per-FedAvg~\citep{NEURIPS2020_24389bfe} \\
    \cmidrule{2-3}
    & Regularization-based pFL & pFedMe~\citep{t2020personalized} and Ditto~\citep{li2021ditto} \\
    \cmidrule{2-3}
    & \multirow{3}{*}{Personalized-aggregation-based pFL} & APFL~\citep{deng2020adaptive}, FedFomo~\citep{zhang2020personalized}, FedAMP~\citep{huang2021personalized}, FedPHP~\citep{li2021fedphp}, APPLE~\citep{ijcai2022p301}, and FedALA~\citep{zhang2023fedala} \\
    \cmidrule{2-3}
    & \multirow{5}{*}{Model-splitting-based pFL} & FedPer~\citep{arivazhagan2019federated}, LG-FedAvg~\citep{liang2020think}, FedRep~\citep{collins2021exploiting}, FedRoD~\citep{chen2021bridging}, FedBABU~\citep{ohfedbabu}, FedGC~\citep{niu2022federated}, FedCP~\citep{Zhang2023fedcp}, GPFL~\citep{zhang2023gpfl}, FedGH~\citep{yi2023fedgh}, DBE~\citep{zhang2023eliminating}, FedCAC~\citep{wu2023bold}, and PFL-DA~\citep{shi2023personalized} \\
    \cmidrule{2-3}
    & \multirow{3}{*}{Knowledge-distillation-based pFL} & FedDistill~\citep{seo202216}, FML~\citep{shen2020federated}, FedKD~\citep{wu2022communication}, FedProto~\citep{tan2022fedproto}, FedPCL~\citep{tan2022federated}, and FedPAC~\citep{xu2022personalized} \\    
    \cmidrule{2-3}
    & Other pFL & FedMTL~\citep{seo202216} and FedBN~\citep{li2020fedbn} \\  
    \bottomrule
    \end{tabular}}
    \label{tab:algo}
\end{table}

\noindent\textbf{Scenarios and Datasets.} 
%According to the popular survey~\citep{li2022federated}, there are mainly three kinds of statistical heterogeneity, including the label distribution skew, the feature distribution skew, and the quantity skew. 
%we mix the quantity skew into the label distribution skew and the feature distribution skew, then we call these 
In \pfllib, we first consider two types of scenarios for data heterogeneity: \textit{label skew} and \textit{feature shift}, where different client datasets differ in label categories and feature categories, respectively~\citep{zhang2023gpfl}. Both CV and NLP classification tasks are considered in these two scenarios. Moreover, we also introduce a \textit{real world} scenario with naturally collected datasets from distributed sensors (HAR~\citep{anguita2012human} and PAMAP2~\citep{reiss2012introducing} for SSP tasks), hospitals (Camelyon17~\citep{koh2021wilds} for CV tasks), and camera traps (iWildCam~\citep{koh2021wilds} for CV tasks) to evaluate algorithms' performance in realistic scenarios. 
% Specifically, we assign the data collected from each single entity (\eg, a sensor or hospital) to each client.

\noindent\textbf{Privacy Evaluation.} We implement the popular Deep Leakage from Gradients (DLG) attack~\citep{zhu2019deep} and the Peak Signal-to-Noise Ratio (PSNR) metric~\citep{wufedcg} to evaluate the privacy-preserving abilities of existing tFL/pFL algorithms. 

\noindent\textbf{Easy to Use and Extend.} 
In \pfllib, algorithms are implemented by \textbf{three} critical files: \linecode{serverX.py} for server creation, \linecode{clientX.py} for client creation, and \linecode{main.py} for hyperparameter configuration. Here, ``\linecode{X}'' represents some algorithm name. We can simply create a new algorithm \linecode{Y} by only adding specific features to \linecode{serverY.py} and \linecode{clientY.py} and while utilizing the core APIs in \linecode{serverbase.py} and \linecode{clientbase.py}. 
To create a scenario, users simply need \textbf{one} command line and run the evaluation with another \textbf{one} command line. 
% \yang{add exactly what are these command lines. maybe use a figure.}
We show an example of using FedALA on MNIST~\citep{lecun1998gradient}:
\begin{lstlisting}[language=Bash]
# generate a practical non-iid and unbalanced scenario using MNIST
python generate_MNIST.py noniid - dir # in ./dataset
# evaluate the FedALA algorithm using a CNN with default hyperparameters
python main.py -data MNIST -m CNN -algo FedALA -gr 2000 -did 0 # in ./system
\end{lstlisting}

\noindent\textbf{Impacts.} 
Our \pfllib is active and popular in the pFL community, as shown by the increasing number of GitHub stars, forks, and active discussions. Due to its simplicity and extensibility, numerous new platforms and projects have been built upon it, such as the FL-bench\footnote{\url{https://github.com/KarhouTam/FL-bench/tree/c11efc286dab4565245da34d7300d5bb07b87a0a}}, the HtFLlib\footnote{\url{https://github.com/TsingZ0/HtFLlib}}, and the FL-IoT\footnote{\url{https://github.com/TsingZ0/FL-IoT}}. Besides, the experiments of several latest SOTA methods~\citep{zhang2023fedala, Zhang2023fedcp, zhang2023gpfl, zhang2023eliminating, zhang2024fedtgp, zhang2024upload, zhang2024fedl2g} are also conducted using our \pfllib. 

\noindent\textbf{Benchmark.} 
Due to limited space here, we only evaluate 20 algorithms in two \textit{label skew} scenarios following the default settings\footnote{Due to frequent updates, some default settings and codes for scenario creation may change in \pfllib.} of GPFL~\citep{zhang2023gpfl}. Please refer to our official website\footnote{\url{https://www.pfllib.com/}} for more documents, details, and results. In \Cref{tab:label_skew}, we use the 4-layer CNN~\citep{mcmahan2017communication} for CV tasks on Fashion-MNIST (FMNIST)~\citep{xiao2017fashion}, Cifar100, and Tiny-ImageNet~\citep{chrabaszcz2017downsampled} (TINY for short) datasets and use the fastText~\citep{joulinetal2017bag} for NLP tasks on AG News~\citep{zhang2015character} dataset. We also use ResNet-18~\citep{he2016deep} on Tiny-ImageNet and denote it TINY*.

\begin{table*}[ht]
  \centering
  \caption{The test accuracy (\%) on the CV and NLP tasks in \textit{label skew} settings. }
  \resizebox{\linewidth}{!}{
    \begin{tabular}{l|ccc|cccccc}
    \toprule
    \textbf{Settings} & \multicolumn{3}{c|}{\textbf{Pathological \textit{Label Skew} Setting}} & \multicolumn{5}{c}{\textbf{Practical \textit{Label Skew} Setting}} \\
    \midrule
     & FMNIST & Cifar100 & TINY & FMNIST & Cifar100 & TINY & TINY* & AG News\\
    \midrule
    \textbf{FedAvg} & 80.41$\pm$0.08 & 25.98$\pm$0.13 & 14.20$\pm$0.47 & 85.85$\pm$0.19 & 31.89$\pm$0.47 & 19.46$\pm$0.20 & 19.45$\pm$0.13 & 87.12$\pm$0.19 \\
    \textbf{FedProx} & 78.08$\pm$0.15 & 25.94$\pm$0.16 & 13.85$\pm$0.25 & 85.63$\pm$0.57 & 31.99$\pm$0.41 & 19.37$\pm$0.22 & 19.27$\pm$0.23 & 87.21$\pm$0.13 \\
    % \textbf{MOON} & 70.67$\pm$0.38  & 25.65$\pm$0.21 & 14.93$\pm$0.42 & 83.69$\pm$0.52 & 32.37$\pm$0.48 & 19.68$\pm$0.20 & 19.02$\pm$0.21 & 84.14$\pm$0.17\\
    \textbf{FedGen}& 79.76$\pm$0.60 & 20.80$\pm$1.00 &13.82$\pm$0.09 & 84.90$\pm$0.31 & 30.96$\pm$0.54 & 19.39$\pm$0.18 & 18.53$\pm$0.32 & 89.86$\pm$0.83\\
    \midrule
    \textbf{Per-FedAvg} & 99.18$\pm$0.08 & 56.80$\pm$0.26 & 28.06$\pm$0.40 & 95.10$\pm$0.10 & 44.28$\pm$0.33 & 25.07$\pm$0.07 & 21.81$\pm$0.54 & 87.08$\pm$0.26 \\
    \midrule
    \textbf{pFedMe} & 99.35$\pm$0.14 & 58.20$\pm$0.14 & 27.71$\pm$0.40 & 97.25$\pm$0.17 & 47.34$\pm$0.46 & 26.93$\pm$0.19 & 33.44$\pm$0.33 & 87.08$\pm$0.18 \\
    \textbf{Ditto} & 99.44$\pm$0.06 & 67.23$\pm$0.07 & 39.90$\pm$0.42 & 97.47$\pm$0.04 & 52.87$\pm$0.64 & 32.15$\pm$0.04 & 35.92$\pm$0.43 & 91.89$\pm$0.17 \\
    \midrule
    \textbf{APFL} & 99.41$\pm$0.02 & 64.26$\pm$0.13 & 36.47$\pm$0.44 & 97.25$\pm$0.08 & 46.74$\pm$0.60 & 34.86$\pm$0.43 & 35.81$\pm$0.37 & 89.37$\pm$0.86\\
    \textbf{FedFomo} & 99.46$\pm$0.01  & 62.49$\pm$0.22 & 36.55$\pm$0.50 & 97.21$\pm$0.02 & 45.39$\pm$0.45 & 26.33$\pm$0.22 & 26.84$\pm$0.11 & 91.20$\pm$0.18 \\
    \textbf{FedAMP} &99.42$\pm$0.03  & 64.34$\pm$0.37 &  36.12$\pm$0.30 &97.20$\pm$0.06 & 47.69$\pm$0.49 & 27.99$\pm$0.11 & 29.11$\pm$0.15 & 83.35$\pm$0.05 \\
    \textbf{APPLE} & 99.30$\pm$0.01 & 65.80$\pm$0.08 &  36.22$\pm$0.40 & 97.06$\pm$0.07 & 53.22$\pm$0.20 & 35.04$\pm$0.47 & 39.93$\pm$0.52 & 84.10$\pm$0.18 \\
    \textbf{FedALA} & 99.57$\pm$0.01 & 67.83$\pm$0.06 & 40.31$\pm$0.30 & 97.66$\pm$0.02 & 55.92$\pm$0.03 & 40.54$\pm$0.02 & 41.94$\pm$0.02 & 92.45$\pm$0.10 \\
    \midrule
    \textbf{FedPer} & 99.47$\pm$0.03 & 63.53$\pm$0.21 & 39.80$\pm$0.39 & 97.44$\pm$0.06 & 49.63$\pm$0.54 & 33.84$\pm$0.34 & 38.45$\pm$0.85 & 91.85$\pm$0.24 \\
    \textbf{FedRep} & 99.56$\pm$0.03 & 67.56$\pm$0.31 & 40.85$\pm$0.37 & 97.56$\pm$0.04 & 52.39$\pm$0.35 & 37.27$\pm$0.20 & 39.95$\pm$0.61 & 92.25$\pm$0.20 \\
    \textbf{FedRoD} & 99.52$\pm$0.05 & 62.30$\pm$0.02 & 37.95$\pm$0.22 & 97.52$\pm$0.04 & 50.94$\pm$0.11 & 36.43$\pm$0.05 & 37.99$\pm$0.26 & 92.16$\pm$0.12 \\
    \textbf{FedBABU} & 99.41$\pm$0.05 & 66.85$\pm$0.07 & 40.72$\pm$0.64 & 97.46$\pm$0.07 & 55.02$\pm$0.33 & 36.82$\pm$0.45 & 34.50$\pm$0.62 & 95.86$\pm$0.41\\
    \textbf{FedCP} & 99.66$\pm$0.04 & 71.80$\pm$0.16 & 44.52$\pm$0.22 & \textbf{97.89$\pm$0.05} & 59.56$\pm$0.08 & \textbf{43.49$\pm$0.04} & \textbf{44.18$\pm$0.21} & 92.89$\pm$0.10 \\
    \textbf{GPFL} & \textbf{99.85$\pm$0.08} & 71.78$\pm$0.26 & \textbf{44.58$\pm$0.06} & 97.81$\pm$0.09 & 61.86$\pm$0.31 & 43.37$\pm$0.53 & 43.70$\pm$0.44 & \textbf{97.97$\pm$0.14} \\
    \textbf{FedDBE} & 99.74$\pm$0.04 & \textbf{73.38$\pm$0.18} & 42.89$\pm$0.29 & 97.69$\pm$0.05 & \textbf{64.39$\pm$0.27} & 43.32$\pm$0.37 & 42.98$\pm$0.52 & 96.87$\pm$0.18 \\
    \midrule
    \textbf{FedDistill} & 99.51$\pm$0.03& 66.78$\pm$0.15 & 37.21$\pm$0.25 & 97.43$\pm$0.04 & 49.93$\pm$0.23 & 30.02$\pm$0.09 & 29.88$\pm$0.41 & 85.76$\pm$0.09 \\
    \textbf{FedProto} & 99.49$\pm$0.04 & 69.18$\pm$0.03 & 36.78$\pm$0.07 & 97.40$\pm$0.02 & 52.70$\pm$0.33 & 31.21$\pm$0.16 & 26.38$\pm$0.40 & 96.34$\pm$0.58 \\
    \bottomrule
    \end{tabular}}
  \label{tab:label_skew}
\end{table*}

\vspace{-5mm}

\section{Conclusion}

To support the rapidly evolving pFL research community, we built \pfllib, a beginner-friendly library that includes 37 cutting-edge tFL / pFL algorithms. Besides, we also built a benchmark platform in \pfllib with comprehensive features, datasets, and scenarios. 
% We are committed to the ongoing development of \pfllib, expanding its range of tFL/pFL algorithms, scenarios, datasets, \etc.

% \acks{All acknowledgements go at the end of the paper before appendices and references.
% Moreover, you are required to declare funding (financial activities supporting the
% submitted work) and competing interests (related financial activities outside the submitted work).
% More information about this disclosure can be found on the JMLR website.}
\acks{This work was supported by the National Key R\&D Program of China under Grant No.2022ZD0160504, the Program of Technology Innovation of the Science and Technology Commission of Shanghai Municipality (Granted No. 21511104700), the Interdisciplinary Program of Shanghai Jiao Tong University (project number YG2024QNB05), Tsinghua University(AIR)-Asiainfo Technologies (China) Inc. Joint Research Center, and the National Natural Science Foundation of China (62472284).}

% Manual newpage inserted to improve layout of sample file - not
% needed in general before appendices/bibliography.

\vskip 0.2in
\bibliography{sample}

\end{document}